### v1

# Real-Time Position-Aware View Synthesis from Single-View Input

Abstract.
Recent advancements in view synthesis have significantly enhanced immersive experiences across various computer graphics and multimedia applications, including telepresence, and entertainment. By enabling the generation of new perspectives from a single input view, view synthesis allows users to better perceive and interact with their environment. However, many state-of-the-art methods, while achieving high visual quality, face limitations in real-time performance, which makes them less suitable for live applications where low latency is critical. In this paper, we present a lightweight, position-aware network designed for real-time view synthesis from a single input image and a target camera pose. The proposed framework consists of a Position Aware Embedding, modeled with a multi-layer perceptron, which efficiently maps positional information from the target pose to generate high dimensional feature maps. These feature maps, along with the input image, are fed into a Rendering Network that merges features from dual encoder branches to resolve both high level semantics and low level details, producing a realistic new view of the scene. Experimental results demonstrate that our method achieves superior efficiency and visual quality compared to existing approaches, particularly in handling complex translational movements without explicit geometric operations like warping. This work marks a step toward enabling real-time view synthesis from a single image for live and interactive applications.
1. Introduction
Telepresence systems (Dima and Sjöström, 2021) enable users to interact with remote environments in real time, offering a wide range of applications, particularly in sectors such as heavy machinery, where remote operation from a safe distance is crucial. These systems have the potential to facilitate seamless natural interaction with remote environments. Current telepresence technologies typically rely on head-mounted displays (HMDs), high-speed networks, and consumer or industrial-grade cameras to deliver live video feeds of the environment. While 2D video feeds provide some pictorial information, the lack of depth perception significantly limits situational awareness, which is vital for effective remote operation.
The challenge to have an understanding of a remote site’s structure may be address by enabling motion parallax for the user. I.e., the user moves the head forth and back to get a notion of the scene depth. This requires access to views from different perspectives, where camera positions and poses need to be continuous for a smooth transition and a high quality experience. Therefore, it is imperative to render continuous perspective views accurately and in real time for applications such as augmented reality (AR), virtual reality (VR), and telepresence systems.
The synthesis of views from perspectives continuous in camera position and pose requires sufficient information of the scene. Early methods employed view interpolations and depth-image-based rendering, but both these require multiple capture devices resulting in calibration issues and high communication capacity requirements. Therefore, generating new views from a single input image and a specified camera pose has emerged as a key task in rendering. A variety of approaches, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021) and more recent transformer-based architectures like NViST (Jang and Agapito, 2024), have been proposed to tackle this problem. These methods leverage both synthetic and real-world datasets to produce high-quality novel views. However, most state-of-the-art methods either require multi-view inputs or are computationally intensive, making them unsuitable for real-time applications where low latency is critical.
In this work, we build upon the growing body of research on single-image-based view synthesis. We introduce a learning based position embedding method combined with a rendering network that efficiently synthesizes high-quality views of arbitrary camera location and pose using fewer computational resources. Our method maintains high fidelity even when handling large translational shifts, making it ideal for real-time applications and practical deployment on consumer devices. To train and test our method we also create our own synthetic dataset using Blender.
Unlike other methods, such as NViST, which employs transformer architectures, or SE3D (Koh et al., 2023), which utilizes large-scale generative-adversarial networks (GANs) to address occlusions and complex scenes, our network is designed to be lightweight and computationally efficient. Despite its smaller footprint, it excels in handling scenarios where translational movements dominate, enabling real-time inference. In summary, our main contributions are as follows:
A lightweight network with a positional embedding module that can generate novel views within a given baseline without requiring explicit operations.
A positional encoding scheme similar to NeRF, which is combined with an multi-layer perceptron (MLP), to create a more generalizable framework capable of handling different scenes.
Our method achieves real-time performance, with inference times of approximately 8.6 milliseconds, corresponding to over 100 frames per second (FPS) for images at a resolution of 512x512 pixels.
The remainder of this paper is organized as follows: Section 2 reviews state-of-the-art methods aimed at view synthesis. Section 3, and 4 describes the proposed approach for small baseline and large baseline rendering. Section 5 provides a comparative analysis of the results, including quality and speed evaluations. Finally, conclusions are summarized in Section 7.
2. Background
View synthesis has recently become a pivotal research area in immersive telepresence, aiming to generate unseen perspectives of a remote scene from a limited set of input images. Traditional approaches, such as structure-from-motion (SfM) and multi-view stereo (MVS), rely on geometric consistency between input views to reconstruct novel viewpoints (Seitz and Dyer, 1999). While these methods have shown effectiveness in controlled environments, their performance deteriorates in complex real-world scenarios and often necessitates multiple input views to achieve high-quality results. In light of these limitations, more recent methods focus on synthesizing novel views from a sparse or single image. This section outlines these advancements and their respective contributions to the field of view synthesis.
2.1. Scene Regression Approaches
Recent advancements in view synthesis have centered around methods that aim to represent a scene from sparse image inputs, particularly those leveraging neural implicit representations such as NeRF (Mildenhall et al., 2021). A NeRF encodes a scene as a continuous volumetric function, allowing novel view generation through volumetric rendering. Despite its capacity to produce highly photorealistic images, NeRF suffers from computational inefficiencies, particularly in terms of inference time. Several techniques have been proposed to mitigate these limitations, such as hash encoding in InstantNGP (Müller et al., 2022) and multi-scale representations in Mip-NeRF (Barron et al., 2021). While these improvements accelerate rendering, they retain significant reliance on MLPs, which remain computationally expensive.
To address these concerns, methods like Plenoxels (Fridovich-Keil et al., 2022) have emerged, leveraging spherical harmonics to represent a scene without the need for deep neural networks. Though Plenoxels achieves faster optimization than many NeRF-based methods, it still falls short in terms of memory efficiency. A more recent innovation, Gaussian Splatting (Kerbl et al., 2023), improves memory and rendering efficiency by employing high-dimensional Gaussian kernels to represent the scene. However, its optimal performance still depends on the availability of multiple input views.
2.2. Light Field Reconstruction and Rendering
Light field (LF) representation encodes both the spatial and angular information of a scene, often described by an HxWxNxN structure, where HxW refers to the spatial resolution and NxN denotes the angular resolution (Levoy and Hanrahan, 1996). This representation facilitates tasks like viewpoint manipulation (Zhou et al., 2020). LF reconstruction methods aim to generate a dense NxN light field from a single input view or a few input views, eliminating the need for complex multi-camera setups.
A common LF reconstruction approach utilizes convolutional neural networks (CNNs) to estimate scene depth (Srinivasan et al., 2017), followed by warping operators to generate target views. To address warping artifacts, subsequent refinement networks have been developed (Zhou et al., 2020, 2021). Other methods exploit depth estimation or optical flow (Cun et al., 2019), appearance flow (Ivan et al., 2019), or a combination of both (Bae et al., 2021) to synthesize the light field.
In more recent developments, GANs (Chen et al., 2020; Chandramouli et al., 2020) and auto-encoders (Han and Xiang, 2022) have demonstrated effectiveness in reconstructing dense light fields from a single image input. Additionally, LFSphereNet (Gond et al., 2023) extends LF reconstruction to spherical light fields using a single omnidirectional image. Furthermore, to generate the novel views from the LFs, the algorithms like dynamically reparameterized LFs (Isaksen et al., 2000) and per-view disk-based blending (Overbeck et al., 2018) have to be utilized which have higher memory footprint, making the whole process unsuitable for dynamic environments and low memory systems.
Our proposed method, PLFNet (described in Sec 3), contributes to LF reconstruction and rendering by directly synthesizing views within a local LF volume, thus eliminating the need of additional rendering algorithms.
2.3. Learning-Based Approaches
Learning-based methods seek to overcome the limitations of both scene regression and light field rendering approaches, which include constraints such as low baselines, sparse image input requirements, and slow rendering times. Several models, including PixelNeRF (Yu et al., 2021) and VisionNeRF (Lin et al., 2023), have been developed to generalize NeRF to novel scenes by conditioning the network on image features. Despite their promise, these methods often face challenges such as scale ambiguity and difficulty in handling large translational movements.
To address these challenges, the diffusion model based approach in 3DiM (Watson et al., 2022) employs pose guided diffusion to improve performance, particularly in scenarios with large translational movement. Fine-tuning diffusion models on large-scale synthetic datasets (Deitke et al., 2023) has also shown potential (Liu et al., 2023, 2024), although these methods are hindered by slow inference times.
An alternative strategy involves using multiplane image (MPI) representations, which approximate a scene using a series of parallel planes at fixed depths, each containing an image and alpha map. This representation allows novel view generation based on a fixed (Zhou et al., 2018; Flynn et al., 2019; Tucker and Snavely, 2020; Li and Kalantari, 2020) or variable (Li and Kalantari, 2020) number of planes. While MPI-based methods achieve superior visual quality, their computational demands remain high, especially during MPI layer prediction. Large-scale GANs (Koh et al., 2023) have also been explored for their ability to render fast and handle large baselines, though they introduce significant memory overhead, limiting their practicality.
More recent innovations, such as NViST (Jang and Agapito, 2024), incorporate vision transformers (ViTs) (Dosovitskiy, 2020) into view synthesis architectures, demonstrating strong generalization capabilities in real-world settings. By leveraging masked autoencoders (MAEs) and adaptive layer normalization conditioned on camera parameters, NViST delivers high-quality results, even for out-of-distribution scenes.
Our proposed method, PLFNet+ (described in Sec 4), contributes to the learning-based category by extending the views from a single input image while diverging from diffusion and transformer-based methods. By utilizing a lightweight design, PLFNet+ reduces computational overhead, enabling real-time rendering. This efficiency makes it particularly suited for dynamic environments where fast processing is crucial, addressing key challenges in the field of view synthesis.
3. Method: Light Field Reconstruction - PLFNet
We first define a Position Aware Light Field Reconstruction network called ”PLFNet” which aims to reconstruct the LF from a single input image by iterative reconstructing individual sub-aperture images. The source image serves as the basis for reconstructing the surrounding LF , where x represents the angular coordinate , and signifies the center of the LF. Each sub-aperture image at location , specified as , undergoes reconstruction via:
This two stage pipeline has been shown in Figure 1
3.1. Positional Embedding
We define the target position by the coordinate vector , where , and represent translations associated with angular coordinates. Given limited baseline shifts of LF reconstruction, directly using normalized values of can suffice for training the network on discrete locations but querying for any continuous values of .
Coordinate Normalization: The target coordinates are normalized into to ensure consistency across varying ranges of , . The normalization process is given by
This normalization step ensures that the coordinates lie within a fixed range, which is crucial for subsequent encoding.
Input Channel: To incorporate the normalized coordinates into the network, the values of are expanded into a tensor , where the first channel holds the values, and the second channel holds the values
This tensor has the same spatial resolution as the input image, ensuring that positional information is aligned with the image features spatially.
3.2. Rendering Network
The role of the rendering network in PLFNet is to synthesize a novel target image from the input image and the positional embedding , obtained from the positional embedding stage. Our network architecture is inspired by LFSphereNet (Gond et al., 2023) and consists of two distinct encoder branches followed by a decoder network. These branches work in tandem to extract and merge spatial and positional features, ultimately generating high quality novel views.
The rendering network operates as follows: the first encoder processes the input image to extract image-based features, while the second encoder processes both the input image and the positional feature map to capture position-dependent information. These features are subsequently merged in the decoder, which refines and upsamples the combined feature maps to generate the final output , a novel view at the target position.
3.2.1. Encoder I - Pre-trained Image Feature Extractor
The first encoder, , is a pre-trained image feature extractor that processes the input image . Following the approach used in ResNet-152 (He et al., 2016), we utilize the initial layers of ResNet to extract lower-level feature maps that are rich in texture and local detail. Specifically, extracts a set of feature maps , where and represent the spatial dimensions of the downsampled feature maps. These feature maps are then expanded to through additional convolution layers to increase the depth of the feature space which can be described as
The extracted features capture essential image content but lack positional awareness, which is handled by the second encoder.
3.2.2. Encoder II - Position-Aware Feature Extractor
The second encoder, , integrates both the input image and the positional embedding to adaptively capture position-specific features. Unlike the first encoder, which only processes the image, focuses on refining the feature maps based on the target position.
We first concatenate the input image and the positional embedding along the channel dimension, resulting in a stacked input tensor
where is the number of channels in . In our case, for LF Reconstruction.
This stacked tensor is passed through a series of convolutional layers that preserve the spatial dimensions, ensuring pixel-level positional information is retained throughout the encoding process, as demonstrated by LFSphereNet (Gond et al., 2023). The encoder produces a feature map , where is the number of feature channels, and skip connections are generated from intermediate layers for later use in the decoder.
Formally, this can be expressed as
The skip connections, denoted as , are collected from intermediate layers, where indexes the corresponding layer, and are used to enhance spatial detail during the decoding stage.
3.2.3. View Decoder
The decoder, , is responsible for merging the feature maps from both encoder branches and reconstructing the target image at the desired target position. The concatenated feature maps and are first combined along the channel dimension
This concatenated feature map is then passed through a series of transposed convolution layers to progressively upsample the spatial resolution, ultimately matching the resolution of the input image. The decoder uses a combination of transposed convolutions and bilinear upsampling, preserving spatial coherence in the generated image. Skip connections from the second encoder are concatenated into the upsampled feature maps at various stages to retain high-frequency details. The skip connections ensure that fine-grained positional and image details are retained, resulting in more consistent and accurate novel views. The overall process can be expressed as
The final output of the decoder is the synthesized target image , with spatial dimensions .
3.3. Training Loss
To optimize our PLFNet, we employ a combination of the L1 loss and the Multi-Scale Structural Similarity (MS-SSIM) (Wang et al., 2003) loss. This combination leverages the complementary strengths of each loss function. The L1 loss minimizes pixel-wise differences between the predicted sub-aperture images and the ground truth images , which helps to maintain accurate reconstruction of intensity values. The MS-SSIM loss, on the other hand, is a perceptually motivated metric that focuses on preserving structural information and texture across multiple scales, which correlates better with human visual perception as shown in image restoration task (Zhao et al., 2016).
The total loss function is defined as a weighted sum of the L1 and MS-SSIM losses
where is a weighting factor that balances the contribution of the two losses. This combined loss function enables the network to generate high-quality light field reconstructions that are both pixel-accurate and perceptually consistent.
3.4. Implementation: PLFNet
The input of PLFNet is a single RGB image and the target position . Here and which are similar to the spatial resolution of input image . We set because the datasets (Srinivasan et al., 2017; Raj et al., 2016; Rerabek and Ebrahimi, 2016) are organized as a 2D grid, with only two angular coordinates considered. PLFNet was implemented in PyTorch and trained with the following hyperparameters: batch size of 24, learning rate of 0.003, adam optimizer, MS-SSIM and L1 loss. After each 30 epochs, the learning rate was decreased by a factor of 0.2. The network was trained for 150 epochs on 2 GPUs (Nvidia A40) within a computing cluster with Intel Xeon Gold 6338 CPUs. The training duration was 48 hrs. For the dataset split, 75% of the data was used for training, 12.5% for validation and 12.5% for testing.
4. Method: View Synthesis - PLFNet+
Our approach for view synthesis builds upon PLFNet by extending the positional embedding to handle wider baselines with translation along the -axis. Here, represents the synthesized image corresponding to the target viewpoint. Our approach consists of two key stages as shown in Figure 2: first, the target position is embedded into a high-dimensional space through positional embedding (discussed in Section 4.1), and second, novel views are rendered from these positions, which is detailed in Section 4.2.
4.1. Extended Positional Embedding
We define the target position by the coordinate vector , where , , and represent translations from the input image . As the baseline increases and the search space extends along the axis, it becomes necessary to project the target coordinates into a higher-dimensional space to accurately capture positional information. To address this, we employ a positional embedding strategy that combines positional encoding, inspired by NeRF (Mildenhall et al., 2021), along with a MLP for enhanced representation of the target coordinate .
Coordinate Normalization: Similar to the normalization strategy in Section 3.1, we convert the target coordinates into their normalized form .
Positional Encoding: To project the normalized coordinates into a higher-dimensional space, we apply positional encoding. This involves concatenating multiple sine and cosine functions of varying frequencies
where is the normalized 3D coordinate vector, represents the encoding depth, and . The parameter is a scaling constant based on the domain of . For our experiments, we use , therefore yields an output of size dimensions, which provides enhanced positional awareness (see Table 6 for ablation study results).
Learning Based Reprojection: Once the positional encoding has been computed, we further enhance the representation by passing it through an MLP. This MLP maps the low-dimensional encoded vector into a higher-dimensional space, and it has proven to work well with scene regression task as shown in NeRF (Mildenhall et al., 2021). This process allows for a more expressive and flexible feature space, and our approach is similar to the adaptive layer normalization techniques used in NViST (Jang and Agapito, 2024).
Let be the output of the positional encoding, where is the dimensionality of the encoding. The MLP projects this vector into a high-dimensional space , where is chosen based on the desired resolution. Mathematically, the projection is represented as
After this, the vector is reshaped into a feature map of size , where . This reshaping operation can be expressed using the inverse vectorization notation as
Finally, the feature map is upsampled to match the spatial resolution of the input image, denoted by . This upsampling is performed using a scaling factor , as follows
The final upsampled representation retains the positional information of the target coordinate and is subsequently used by the rendering network to generate the novel view.
4.2. Rendering Network
The rendering network remains similar to Section 3.2. However, in the case of view synthesis, the positional embedding is denoted by a single channel, simplifying the input to the second encoder as
The output of the rendering network is the target image .
4.3. Training Loss
For the view synthesis task, we extend the loss function previously defined in Section 3.3 to better capture fine-grained details, especially with larger baselines. We maintain the same combination of L1 loss and MS-SSIM loss, which focuses on minimizing pixel-wise differences and preserving structural consistency, respectively. However, to further enhance the quality of the synthesized views, particularly in the frequency domain, we incorporate the Focal Frequency Loss (FFL) (Jiang et al., 2021) into the total loss function.
The FFL addresses the spectral bias in neural networks by adaptively focusing on the frequency components that are hardest to synthesize. This allows the model to improve the reconstruction of high-frequency details, which are critical for realistic novel view generation.
The FFL is defined as
where and represent the real and predicted frequency components at position in the Fourier domain, and is a dynamic spectrum weight that emphasizes hard-to-synthesize frequencies.
Thus, the total loss function for PLFNet+ is updated to
where is the Focal Frequency Loss, and is a weight balancing its contribution relative to the L1 and MS-SSIM losses.
This augmented loss function allows PLFNet+ to generate sharper, more realistic novel views by not only optimizing for spatial accuracy but also addressing frequency domain discrepancies.
4.4. Implementation: PLFNet+
We keep the same implementation as in PLFNet (described in Section 3.4, but set , . We also set as the positional maps is single channel. Due to limitation of training data when using our synthetic dataset we employ multi stage training strategies to reach satisfactory results in training process.
4.4.1. Multi Stage Training
Given the scarcity of large-scale datasets for view synthesis with wide baselines, we adopt a three-stage training approach shown in Figure 3 to progressively enhance the network’s ability to capture image features and handle positional shifts.
Stage 1 serves as a preconditioning step, training the network on a large-scale light field dataset (Srinivasan et al., 2017). While the baseline between the views in each light field is relatively narrow, the dataset’s large scale enables the network to learn both high- and low-level image features, implicitly understanding depth. In this stage, the network is conditioned to synthesize novel views within a constrained light field volume, similar to the reconstruction process described in PLFNet (Section 3). Once the network demonstrates proficiency in this task, we transition to the next stage.
Stage 2 introduces a dataset with a significantly larger baseline: the Spherical Light Field Database (SLFDB) (Zerman et al., 2024), which consists of 20 spherical light fields captured with 60 angular dimensions. From these omnidirectional images (ODIs), we extract planar patches and fine-tune the network with a lower learning rate. This step enables the network to learn to handle wider baseline movements while maintaining positional accuracy. Following this, the network is prepared for the final stage of training, which focuses on the joint optimization of the MLP and rendering network with more random positions.
Stage 3 involves training the network on a synthetic dataset containing random camera viewpoints within an cubical volume. The increase in the spatial search space necessitates the use of the positional embedding module for accurate view synthesis. In this stage, the network is fine-tuned to handle more complex camera translations and larger scene variations, fully utilizing the positional embedding to enhance rendering precision.
5. Experimental Evaluation
In this section, we present a comprehensive evaluation of our proposed PLFNet and PLFNet+ architectures. We begin by assessing the performance of PLFNet+ for view synthesis on datasets with varying baselines and complexities. Following this, we demonstrate the capability of PLFNet in reconstructing full light fields from a single input image. The experiments for view synthesis are detailed in Section 5.1, while light field reconstruction results are discussed in Section 5.2.
5.1. View Synthesis
| Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| SinMPI (Pu et al., 2023) | 17.3295 | 0.4193 | 0.7064 | 0.4911 | 0.2794 | 0.4097 | |
| Blender | TMPI (Khan et al., 2023) | 19.9507 | 0.7160 | 0.8264 | 0.4822 | 0.1458 | 0.1471 |
| 256x256 | AdaMPI (Han et al., 2022) | 24.9252 | 0.8092 | 0.9145 | 0.6720 | 0.0680 | 0.0455 |
| PLFNet+ (Ours) | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 | |
| SinMPI (Pu et al., 2023) | 18.3441 | 0.4599 | 0.6885 | 0.3621 | 0.2251 | 0.3966 | |
| Blender | TMPI (Khan et al., 2023) | 20.0126 | 0.7071 | 0.8070 | 0.3605 | 0.1268 | 0.1640 |
| 512x512 | AdaMPI (Han et al., 2022) | 23.1416 | 0.7510 | 0.8566 | 0.4353 | 0.0751 | 0.0770 |
| PLFNet+ (Ours) | 26.3339 | 0.8334 | 0.9307 | 0.5633 | 0.2191 | 0.1947 | |
| SinMPI (Pu et al., 2023) | 14.5548 | 0.4271 | 0.6621 | 0.3706 | 0.2867 | 0.4024 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 17.1759 | 0.5454 | 0.7314 | 0.3819 | 0.2028 | 0.2415 |
| 256x256 | AdaMPI (Han et al., 2022) | 23.6155 | 0.8263 | 0.9345 | 0.7232 | 0.1055 | 0.0797 |
| PLFNet+ (Ours) | 22.8933 | 0.7608 | 0.9070 | 0.5534 | 0.2378 | 0.2560 | |
| SinMPI (Pu et al., 2023) | 15.7272 | 0.4740 | 0.6712 | 0.3095 | 0.2392 | 0.3775 | |
| COCO (Caesar et al., 2018) | TMPI (Khan et al., 2023) | 18.1375 | 0.5444 | 0.7499 | 0.2906 | 0.1750 | 0.2758 |
| 512x512 | AdaMPI (Han et al., 2022) | 21.5931 | 0.7425 | 0.8925 | 0.5215 | 0.1156 | 0.1559 |
| PLFNet+ (Ours) | 20.2596 | 0.6229 | 0.8001 | 0.3007 | 0.2758 | 0.4226 |
| Method | 256x256 | 512x512 |
| SinMPI (Pu et al., 2023) | 0.0752 / 13 | 0.2368 / 4 |
| TMPI (Khan et al., 2023) | 0.0910 /11 | 0.2640 / 4 |
| AdaMPI (Han et al., 2022) | 0.2261 / 4 | 0.3254 / 3 |
| PLFNet+ (Ours) | 0.0074 / 135 | 0.0086 / 116 |
5.1.1. Datasets
To evaluate the performance of PLFNet+ on view synthesis, we used two datasets. The first is a custom Blender dataset, created specifically for this task, which contains images from different scenes with camera movements constrained within a 20 cm cubical volume. The second dataset is COCO (Caesar et al., 2018), which we use during training in the same manner as AdaMPI (Han et al., 2022), where depth-image-based rendering (DIBR) with hole filling is employed to generate the training data. During evaluation, we take an input image and synthesize a novel image at a specified target location. The synthesized image is then treated as input to generate an image back at the original location. This process enables us to compare the reconstructed image with the ground truth, which is the original input image, allowing for an assessment of rendering quality. Both datasets are evaluated at resolutions of and and a movement within a volume of .
5.1.2. Evaluation Metrics
To objectively evaluate the quality of the synthesized novel views, we employed multiple standard and advanced metrics. These include peak signal-to-noise ratio (PSNR) and structural similarity index measure (SSIM) (Wang et al., 2004), as well as multi-scale structural similarity index measure (MS-SSIM) (Wang et al., 2003), pixel- based visual information fidelity (VIFP) (Sheikh and Bovik, 2006), deep image structure and texture similarity (DISTS) (Ding et al., 2020), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018). These metrics capture not only pixel-level errors but also perceptual differences, providing a holistic view of the synthesis quality.
5.1.3. Results
Quantitative Results: The quantitative results for view synthesis across different datasets and resolutions are presented in Table 1. Our proposed PLFNet+ consistently outperforms AdaMPI (Han et al., 2022) on the Blender dataset and achieves competitive results on the COCO (Caesar et al., 2018) dataset. Although AdaMPI demonstrates slightly better performance in perceptual metrics such as DISTS and LPIPS, our method remains competitive across a range of metrics, especially on PSNR and SSIM.
Interestingly, we observe a slight performance drop in AdaMPI and PLFNet+ when increasing the resolution from to . This highlights the challenges of scaling models without increasing their capacity to handle higher-resolution inputs. In contrast to other methods that explicitly rely on depth estimation using the DPT model, our approach learns scene depth implicitly from RGB input alone, demonstrating the flexibility of our method without the need for pre-computed depth maps.
When evaluating inference time, PLFNet+ excels, particularly at higher resolutions. As shown in Table 2, our method achieves an inference time of 0.0086 seconds for a resolution of , outperforming all other methods by a significant margin. This makes PLFNet+ highly suitable for real-time applications, where rapid generation of novel views is critical, such as in dynamic video streams or interactive virtual environments.
Qualitative Results: Figure 4 provides a visual comparison of the view synthesis results for different methods on the COCO dataset. Each method’s output is accompanied by error maps, allowing for a more detailed analysis of their accuracy.
SinMPI (Pu et al., 2023) struggles with rendering artifacts, particularly in regions with complex textures and geometry, as evidenced by the error maps. These issues are likely a result of its depth estimation and scene representation. TMPI (Khan et al., 2023) also displays notable tiling artifacts due to its approach of partitioning the scene into multiple planes, leading to visible seams and discontinuities.
In contrast, both AdaMPI and our proposed PLFNet+ produce the most visually accurate results, with minimal artifacts. AdaMPI shows slightly better performance in some scenarios, particularly in handling fine details, but our method’s ability to efficiently handle reflections, occlusions, and fine textures stands out. While AdaMPI may have a slight edge in perceptual quality, our method’s superior inference speed makes it the more practical choice for real-time applications, balancing high-quality synthesis with the demands of time-sensitive environments.
5.2. Extending View Synthesis: Light Field Reconstruction
5.2.1. Datasets
For LF reconstruction, we evaluated our method using three publicly available datasets: Lytro Flowers (Srinivasan et al., 2017), Stanford Light Field Archive (Raj et al., 2016), and JPEG-Pleno (Rerabek and Ebrahimi, 2016). These datasets are widely used for benchmarking light field reconstruction algorithms. The extracted light field images had dimensions of 352x512 with a 7x7 angular resolution.
We trained PLFNet on 75% of the Lytro Flowers and Stanford datasets, with the remaining 25% reserved for testing. The JPEG-Pleno dataset, however, was exclusively used for testing to evaluate the generalization capability of our model, as it was not included in the training process.
5.2.2. Evaluation Metrics
We utilize similar evaluation metrics as described in Section 5.1.2. Since the light field reconstruction works in very narrow baseline between each sub-aperture image, and has discrete camera locations, the search space is much smaller therefore we get much better scores in these metrics.
| Dataset | Method | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| Flowers (Srinivasan et al., 2017) | NoisyLFRecon | 39.9500 | 0.9763 | 0.9932 | 0.9285 | 0.0389 | 0.0190 |
| DGLF | 35.6194 | 0.8773 | 0.9301 | 0.6021 | 0.1556 | 0.1431 | |
| DALF | 37.3006 | 0.8941 | 0.9589 | 0.7362 | 0.1033 | 0.0911 | |
| IR-V | 37.9034 | 0.9122 | 0.9645 | 0.7324 | 0.0994 | 0.0707 | |
| LFSphereNet | 41.3719 | 0.9461 | 0.9868 | 0.9060 | 0.0812 | 0.0512 | |
| PLFNet | 37.2606 | 0.9601 | 0.9914 | 0.9062 | 0.0608 | 0.0229 | |
| Stanford (Raj et al., 2016) | NoisyLFRecon | 35.4917 | 0.9572 | 0.9771 | 0.8380 | 0.0718 | 0.0395 |
| DGLF | 35.6509 | 0.9673 | 0.9897 | 0.8655 | 0.0685 | 0.0321 | |
| DALF | 38.2996 | 0.9172 | 0.9750 | 0.7829 | 0.0667 | 0.0402 | |
| IR-V | 39.4909 | 0.9440 | 0.9852 | 0.8358 | 0.0628 | 0.0390 | |
| LFSphereNet | 40.9830 | 0.9488 | 0.9797 | 0.8842 | 0.0556 | 0.0300 | |
| PLFNet | 36.8530 | 0.9601 | 0.9927 | 0.8679 | 0.0553 | 0.0182 | |
| JPEG Pleno (Rerabek and Ebrahimi, 2016) | NoisyLFRecon | 36.6275 | 0.9575 | 0.9687 | 0.8549 | 0.0718 | 0.0435 |
| DGLF | 32.3387 | 0.7207 | 0.8166 | 0.4246 | 0.1556 | 0.1431 | |
| DALF | 35.0285 | 0.8257 | 0.9291 | 0.6462 | 0.1095 | 0.0937 | |
| IR-V | 37.2085 | 0.9122 | 0.9719 | 0.7878 | 0.0894 | 0.0605 | |
| LFSphereNet | 39.2624 | 0.9429 | 0.9879 | 0.9069 | 0.0737 | 0.0425 | |
| PLFNet | 33.8982 | 0.9106 | 0.9838 | 0.8683 | 0.0590 | 0.0261 |
| Method | Purpose | Inference Time (s) |
| NoisyLFRecon | LF reconstruction | 1.1830 |
| DGLF | LF reconstruction | 1.5009 |
| DALF | LF reconstruction | 0.0583 |
| IR-V | LF reconstruction | 0.1429 |
| LFSphereNet | LF reconstruction | 0.0008 |
| PLFNet | Novel view synthesis | 0.0980 |
| NoisyLFRecon | DGLF | DALF | IR-V | LFSphereNet | PLFNet |
| (Zhou et al., 2021) | (Zhou et al., 2020) | (Cun et al., 2019) | (Han and Xiang, 2022) | (Gond et al., 2023) | (Ours) |
5.2.3. Results
Quantitative Results: The quantitative results of our light field reconstruction experiments are presented in Table 3. Our proposed PLFNet demonstrates competitive performance compared to state-of-the-art methods, including NoisyLFRecon (Zhou et al., 2021), DGLF (Zhou et al., 2020), DALF (Cun et al., 2019), IR-V (Han and Xiang, 2022), and LFSphereNet (Gond et al., 2023). All models were trained from scratch using identical datasets for a fair comparison. Publicly available implementations were used for the methods by Zhou et al. (2021, 2020) and Han and Xiang (2022), while DALF (Cun et al., 2019) was re-implemented in PyTorch according to the original paper.
Although DGLF and LFSphereNet generate the full light field in a single pass, both DALF and our PLFNet reconstruct each sub-aperture image iteratively. For the Stanford dataset, which offers fewer training samples compared to Lytro Flowers, LFSphereNet delivered the best results across most metrics, with PLFNet showing competitive performance, particularly excelling in SSIM and MS-SSIM. In terms of computational efficiency, PLFNet demonstrated significant improvements as shown in Table 4. Our PLFNet achieved an inference time of 0.0980 seconds for whole LF reconstruction which included forward passes. However, per view inference on average is around 0.006 seconds highlighting its potential for real-time applications.
Qualitative Results: The qualitative evaluation of light field reconstruction is illustrated in Figure 5. We compare the visual performance of our proposed PLFNet against state-of-the-art methods for light field views, focusing on the top right corner sub-aperture view. The first two rows depict reconstruction results for the JPEG-Pleno (Rerabek and Ebrahimi, 2016) dataset. In these challenging scenarios, PLFNet demonstrates superior reconstruction quality with fewer visible artifacts, as evident in the error maps. Compared to LFSphereNet (Gond et al., 2023), which also shows strong performance, PLFNet exhibits more accurate details and reduced error magnitudes, highlighting its robustness and generalization capabilities. The subsequent rows display results for the Stanford Light Field Archive (Raj et al., 2016) dataset. Here, our method consistently outperforms NoisyLFRecon (Zhou et al., 2021), which utilizes nine input images for reconstruction. Despite this advantage, NoisyLFRecon produces noticeable artifacts and blurring in regions with fine structures. In contrast, PLFNet, which relies on only a single input image, achieves sharper reconstructions with more faithful reproduction of scene geometry and textures. This indicates that PLFNet can effectively capture and synthesize complex light field information even with minimal input data.
Overall, the qualitative results corroborate the quantitative findings, showcasing ability of PLFNet to achieve high-quality reconstructions with low error rates. This performance shows potential for practical applications where high-fidelity light field reconstruction from sparse data is required.
6. Ablation Study
In this section, we present a two-part ablation study to assess the effectiveness of our proposed method, focusing on positional understanding and the impact of different positional embedding techniques. The first part evaluates the performance of PLFNet+ under varying sparse input configurations, demonstrating how the number and location of input images influence view synthesis. The second part investigates the contribution of different positional encoding strategies to the overall reconstruction quality, highlighting the role of each component in our final model.
6.1. Sparse Input Mode
The first part of the ablation study explores how the performance of PLFNet+ is affected by varying the number and position of input images for view synthesis. We evaluate three configurations: (i) a single central image as input, (ii) a single image from a random location within the cubical volume, and (iii) two images from random locations. These setups allow us to understand how well the model generalizes to sparse input scenarios, particularly in cases where inputs are not centrally aligned.
Table 5 presents the quantitative results of this study. As expected, the single central view input yields the best performance across all metrics, with a PSNR of 28.8023 and SSIM of 0.9009, indicating that the model performs best when provided with a centrally located reference image. The random input configurations, whether with one or two images, exhibit slightly reduced performance, suggesting that while the model can generalize to inputs from random locations, the quality of the synthesized view degrades with less structured input data. Nevertheless, even with these random inputs, PLFNet+ achieves acceptable performance, demonstrating its robustness in handling diverse input scenarios.
| Number of Input Images | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| 2 Images - Random Location | 23.0571 | 0.7444 | 0.8462 | 0.4297 | 0.2782 | 0.2510 |
| 1 Image - Random Location | 22.8429 | 0.7243 | 0.8386 | 0.3524 | 0.3373 | 0.3703 |
| 1 Image - Central View | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
6.2. Positional Embedding
In the second part of the ablation study, we analyze the impact of different positional embedding strategies on view synthesis. We compare four configurations: (i) MLP only, (ii) normalization only, (iii) normalization with positional encoding, (iv) normalization with MLP, and (v) the full setup with normalization, positional encoding, and MLP. This analysis aims to determine the contribution of each component in enhancing the model’s understanding of positional information.
The results are summarized in Table 6. The combination of normalization, positional encoding, and MLP consistently outperforms the other configurations, achieving a PSNR of 28.8023 and SSIM of 0.9003. This demonstrates the effectiveness of our final proposed method in capturing positional variations and synthesizing accurate novel views. By contrast, using only normalization or MLP leads to a noticeable drop in performance specially in high frequency regions, underscoring the importance of combining these techniques to achieve optimal results.
The setup with normalization and MLP alone performs better than using MLP in isolation but still falls short of the complete configuration. These findings highlight the critical role of positional encoding and MLP in providing a richer representation of the 3D positional information, which is essential for accurately reconstructing complex novel views.
| Positional Embedding | PSNR | SSIM | MS-SSIM | VIFP | DISTS | LPIPS |
| MLP | 23.2578 | 0.7633 | 0.8686 | 0.4063 | 0.2566 | 0.2726 |
| Normalization | 25.2855 | 0.8376 | 0.9256 | 0.5309 | 0.2132 | 0.1680 |
| Norm+Positional Encoding | 24.1945 | 0.7941 | 0.8962 | 0.4502 | 0.2276 | 0.1946 |
| Norm+MLP | 26.4475 | 0.8508 | 0.9421 | 0.6247 | 0.2119 | 0.1647 |
| Norm+PosEnc+MLP | 28.8023 | 0.9003 | 0.9655 | 0.7516 | 0.1586 | 0.1007 |
7. Conclusion
In this paper we presented a novel, lightweight, position-aware network for real-time view synthesis and light field reconstruction, which leverages a position-aware embedding combined with a rendering network to achieve both efficiency and high visual quality. Through the development of PLFNet and its extension, PLFNet+, our approach demonstrates superior synthesis quality, particularly in handling complex translational movements without relying on computationally intensive operations like warping.
Extensive evaluations reveal that PLFNet+ achieves a strong balance between visual fidelity and inference speed, positioning it as a feasible solution for applications requiring real-time performance, such as telepresence and augmented reality. Additionally, our network architecture shows robustness in various scenarios, from narrow to wide baseline settings, making it adaptable to dynamic and interactive environments.
The promising results from ablation studies underscore the effectiveness of our positional embedding design, particularly the integration of normalization, positional encoding, and MLP components. Overall, our work provides a significant step towards more accessible, high-performance view synthesis frameworks, paving the way for broader application in consumer and industrial devices.
Acknowledgements.
The work was supported by the European Joint Doctoral Programme on Plenoptic Imaging (PLENOPTIMA) through the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie Grant Agreement No. 956770. The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS), partially funded by the Swedish Research Council through grant agreement no. 2022-06725. We thank the High Performance Computing Center North (HPC2N) at Umeå University for providing computational resources and valuable support during test and performance runs.References
- (1)
- Bae et al. (2021) Kyuho Bae, Andre Ivan, Hajime Nagahara, and In Kyu Park. 2021. 5d light field synthesis from a monocular video. In 2020 25th International Conference on Pattern Recognition (ICPR). IEEE, 7157–7164.
- Barron et al. (2021) Jonathan T Barron, Ben Mildenhall, Matthew Tancik, Peter Hedman, Ricardo Martin-Brualla, and Pratul P Srinivasan. 2021. Mip-nerf: A multiscale representation for anti-aliasing neural radiance fields. In Proceedings of the IEEE/CVF international conference on computer vision. 5855–5864.
- Caesar et al. (2018) Holger Caesar, Jasper Uijlings, and Vittorio Ferrari. 2018. Coco-stuff: Thing and stuff classes in context. In Proceedings of the IEEE conference on computer vision and pattern recognition. 1209–1218.
- Chandramouli et al. (2020) Paramanand Chandramouli, Kanchana Vaishnavi Gandikota, Andreas Goerlitz, Andreas Kolb, and Michael Moeller. 2020. A generative model for generic light field reconstruction. IEEE Transactions on Pattern Analysis and Machine Intelligence 44, 4 (2020), 1712–1724.
- Chen et al. (2020) Bin Chen, Lingyan Ruan, and Miu-Ling Lam. 2020. LFGAN: 4D Light Field Synthesis from a Single RGB Image. ACM Trans. Multimedia Comput. Commun. Appl. 16 (2 2020). Issue 1. https://doi.org/10.1145/3366371
- Cun et al. (2019) Xiaodong Cun, Feng Xu, Chi-Man Pun, and Hao Gao. 2019. Depth-Assisted Full Resolution Network for Single Image-Based View Synthesis. IEEE Computer Graphics and Applications 39 (2019), 52–64. Issue 2. https://doi.org/10.1109/MCG.2018.2884188
- Deitke et al. (2023) Matt Deitke, Dustin Schwenk, Jordi Salvador, Luca Weihs, Oscar Michel, Eli VanderBilt, Ludwig Schmidt, Kiana Ehsani, Aniruddha Kembhavi, and Ali Farhadi. 2023. Objaverse: A universe of annotated 3d objects. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 13142–13153.
- Dima and Sjöström (2021) Elijs Dima and Mårten Sjöström. 2021. Camera and Lidar-Based View Generation for Augmented Remote Operation in Mining Applications. IEEE Access 9 (2021), 82199–82212.
- Ding et al. (2020) Keyan Ding, Kede Ma, Shiqi Wang, and Eero P Simoncelli. 2020. Image quality assessment: Unifying structure and texture similarity. IEEE transactions on pattern analysis and machine intelligence 44, 5 (2020), 2567–2581.
- Dosovitskiy (2020) Alexey Dosovitskiy. 2020. An image is worth 16x16 words: Transformers for image recognition at scale. arXiv preprint arXiv:2010.11929 (2020).
- Flynn et al. (2019) John Flynn, Michael Broxton, Paul Debevec, Matthew DuVall, Graham Fyffe, Ryan Overbeck, Noah Snavely, and Richard Tucker. 2019. DeepView: View Synthesis With Learned Gradient Descent. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR).
- Fridovich-Keil et al. (2022) Sara Fridovich-Keil, Alex Yu, Matthew Tancik, Qinhong Chen, Benjamin Recht, and Angjoo Kanazawa. 2022. Plenoxels: Radiance fields without neural networks. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 5501–5510.
- Gond et al. (2023) Manu Gond, Emin Zerman, Sebastian Knorr, and Mårten Sjöström. 2023. LFSphereNet: Real Time Spherical Light Field Reconstruction from a Single Omnidirectional Image. In Proceedings of the 20th ACM SIGGRAPH European Conference on Visual Media Production. 1–10.
- Han and Xiang (2022) Kang Han and Wei Xiang. 2022. Inference-Reconstruction Variational Autoencoder for Light Field Image Reconstruction. IEEE Transactions on Image Processing 31 (2022), 5629–5644. https://doi.org/10.1109/TIP.2022.3197976
- Han et al. (2022) Yuxuan Han, Ruicheng Wang, and Jiaolong Yang. 2022. Single-view view synthesis in the wild with learned adaptive multiplane images. In ACM SIGGRAPH 2022 Conference Proceedings. 1–8.
- He et al. (2016) Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun. 2016. Deep residual learning for image recognition. In Proceedings of the IEEE conference on computer vision and pattern recognition. 770–778.
- Isaksen et al. (2000) Aaron Isaksen, Leonard McMillan, and Steven J Gortler. 2000. Dynamically reparameterized light fields. In Proceedings of the 27th annual conference on Computer graphics and interactive techniques. 297–306.
- Ivan et al. (2019) Andre Ivan, In Kyu Park, et al. 2019. Synthesizing a 4D spatio-angular consistent light field from a single image. arXiv preprint arXiv:1903.12364 (2019).
- Jang and Agapito (2024) Wonbong Jang and Lourdes Agapito. 2024. NViST: In the Wild New View Synthesis from a Single Image with Transformers. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 10181–10193.
- Jiang et al. (2021) Liming Jiang, Bo Dai, Wayne Wu, and Chen Change Loy. 2021. Focal frequency loss for image reconstruction and synthesis. In Proceedings of the IEEE/CVF international conference on computer vision. 13919–13929.
- Kerbl et al. (2023) Bernhard Kerbl, Georgios Kopanas, Thomas Leimkühler, and George Drettakis. 2023. 3D Gaussian Splatting for Real-Time Radiance Field Rendering. ACM Trans. Graph. 42, 4 (2023), 139–1.
- Khan et al. (2023) Numair Khan, Lei Xiao, and Douglas Lanman. 2023. Tiled multiplane images for practical 3d photography. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 10454–10464.
- Koh et al. (2023) Jing Yu Koh, Harsh Agrawal, Dhruv Batra, Richard Tucker, Austin Waters, Honglak Lee, Yinfei Yang, Jason Baldridge, and Peter Anderson. 2023. Simple and effective synthesis of indoor 3d scenes. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 37. 1169–1178.
- Levoy and Hanrahan (1996) Marc Levoy and Pat Hanrahan. 1996. Light field rendering. In Proceedings of the 23rd annual conference on Computer graphics and interactive techniques. 31–42.
- Li and Kalantari (2020) Qinbo Li and Nima Khademi Kalantari. 2020. Synthesizing Light Field from a Single Image with Variable MPI and Two Network Fusion. ACM Trans. Graph. 39 (11 2020). Issue 6. https://doi.org/10.1145/3414685.3417785
- Lin et al. (2023) Kai-En Lin, Yen-Chen Lin, Wei-Sheng Lai, Tsung-Yi Lin, Yi-Chang Shih, and Ravi Ramamoorthi. 2023. Vision transformer for nerf-based view synthesis from a single input image. In Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision. 806–815.
- Liu et al. (2024) Minghua Liu, Chao Xu, Haian Jin, Linghao Chen, Mukund Varma T, Zexiang Xu, and Hao Su. 2024. One-2-3-45: Any single image to 3d mesh in 45 seconds without per-shape optimization. Advances in Neural Information Processing Systems 36 (2024).
- Liu et al. (2023) Ruoshi Liu, Rundi Wu, Basile Van Hoorick, Pavel Tokmakov, Sergey Zakharov, and Carl Vondrick. 2023. Zero-1-to-3: Zero-shot one image to 3d object. In Proceedings of the IEEE/CVF international conference on computer vision. 9298–9309.
- Mildenhall et al. (2021) Ben Mildenhall, Pratul P. Srinivasan, Matthew Tancik, Jonathan T. Barron, Ravi Ramamoorthi, and Ren Ng. 2021. NeRF: Representing Scenes as Neural Radiance Fields for View Synthesis. Commun. ACM 65, 1 (dec 2021), 99–106. https://doi.org/10.1145/3503250
- Müller et al. (2022) Thomas Müller, Alex Evans, Christoph Schied, and Alexander Keller. 2022. Instant neural graphics primitives with a multiresolution hash encoding. ACM transactions on graphics (TOG) 41, 4 (2022), 1–15.
- Overbeck et al. (2018) Ryan S Overbeck, Daniel Erickson, Daniel Evangelakos, Matt Pharr, and Paul Debevec. 2018. A system for acquiring, processing, and rendering panoramic light field stills for virtual reality. ACM Transactions on Graphics (TOG) 37, 6 (2018), 1–15.
- Pu et al. (2023) Guo Pu, Peng-Shuai Wang, and Zhouhui Lian. 2023. SinMPI: Novel View Synthesis from a Single Image with Expanded Multiplane Images. In SIGGRAPH Asia 2023 Conference Papers. 1–10.
- Raj et al. (2016) Abhilash Sunder Raj, Michael Lowney, Raj Shah, and Gordon Wetzstein. 2016. Stanford lytro light field archive.
- Rerabek and Ebrahimi (2016) Martin Rerabek and Touradj Ebrahimi. 2016. New light field image dataset. In 8th International Conference on Quality of Multimedia Experience (QoMEX).
- Seitz and Dyer (1999) Steven M Seitz and Charles R Dyer. 1999. Photorealistic scene reconstruction by voxel coloring. International journal of computer vision 35 (1999), 151–173.
- Sheikh and Bovik (2006) Hamid R Sheikh and Alan C Bovik. 2006. Image information and visual quality. IEEE Transactions on image processing 15, 2 (2006), 430–444.
- Srinivasan et al. (2017) Pratul P Srinivasan, Tongzhou Wang, Ashwin Sreelal, Ravi Ramamoorthi, and Ren Ng. 2017. Learning to Synthesize a 4D RGBD Light Field From a Single Image. Proceedings of the IEEE International Conference on Computer Vision (ICCV). https://openaccess.thecvf.com/content_iccv_2017/html/Srinivasan_Learning_to_Synthesize_ICCV_2017_paper.html
- Tucker and Snavely (2020) Richard Tucker and Noah Snavely. 2020. Single-view view synthesis with multiplane images. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 551–560.
- Wang et al. (2004) Zhou Wang, A.C. Bovik, H.R. Sheikh, and E.P. Simoncelli. 2004. Image quality assessment: from error visibility to structural similarity. IEEE Transactions on Image Processing 13, 4 (2004), 600–612. https://doi.org/10.1109/TIP.2003.819861
- Wang et al. (2003) Zhou Wang, Eero P Simoncelli, and Alan C Bovik. 2003. Multiscale structural similarity for image quality assessment. In The Thrity-Seventh Asilomar Conference on Signals, Systems & Computers, 2003, Vol. 2. Ieee, 1398–1402.
- Watson et al. (2022) Daniel Watson, William Chan, Ricardo Martin-Brualla, Jonathan Ho, Andrea Tagliasacchi, and Mohammad Norouzi. 2022. Novel view synthesis with diffusion models. arXiv preprint arXiv:2210.04628 (2022).
- Yu et al. (2021) Alex Yu, Vickie Ye, Matthew Tancik, and Angjoo Kanazawa. 2021. pixelnerf: Neural radiance fields from one or few images. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition. 4578–4587.
- Zerman et al. (2024) Emin Zerman, Manu Gond, Soheib Takhtardeshir, Roger Olsson, and Mårten Sjöström. 2024. A Spherical Light Field Database for Immersive Telecommunication and Telepresence Applications. In 2024 16th International Conference on Quality of Multimedia Experience (QoMEX). 200–206. https://doi.org/10.1109/QoMEX61742.2024.10598264
- Zhang et al. (2018) Richard Zhang, Phillip Isola, Alexei A Efros, Eli Shechtman, and Oliver Wang. 2018. The Unreasonable Effectiveness of Deep Features as a Perceptual Metric. In CVPR.
- Zhao et al. (2016) Hang Zhao, Orazio Gallo, Iuri Frosio, and Jan Kautz. 2016. Loss functions for image restoration with neural networks. IEEE Transactions on computational imaging 3, 1 (2016), 47–57.
- Zhou et al. (2018) Tinghui Zhou, Richard Tucker, John Flynn, Graham Fyffe, and Noah Snavely. 2018. Stereo Magnification: Learning View Synthesis Using Multiplane Images. ACM Trans. Graph. 37, 4, Article 65 (jul 2018), 12 pages. https://doi.org/10.1145/3197517.3201323
- Zhou et al. (2020) Wenhui Zhou, Gaomin Liu, Jiangwei Shi, Hua Zhang, and Guojun Dai. 2020. Depth-guided view synthesis for light field reconstruction from a single image. Image and Vision Computing 95 (2020), 103874. https://doi.org/10.1016/j.imavis.2020.103874
- Zhou et al. (2021) Wenhui Zhou, Jiangwei Shi, Yongjie Hong, Lili Lin, and Ercan Engin Kuruoglu. 2021. Robust dense light field reconstruction from sparse noisy sampling. Signal Processing 186 (9 2021). https://doi.org/10.1016/j.sigpro.2021.108121